OpenStudy (anonymous):

Given the functions f(x) = 3x - 1 and g(x) = 3x + 4, which operation results in the smallest coefficient on the x term?

OpenStudy (anonymous):

Addition Subtraction Multiplication Two operations result in the same coefficient

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

@Cameron23

OpenStudy (anonymous):

3x - 3x = 0 3x + 3x = 6x 3x * 3x = 9x Which is smallest?

OpenStudy (anonymous):

1st one